OpenStudy (anonymous):

MEDAL REWARDED (45x^3/7)(3x^3/2)

OpenStudy (anonymous):

135x^6/14

OpenStudy (anonymous):

its 0

OpenStudy (anonymous):

nope its 135x^1/2 haha

OpenStudy (anonymous):

\[\left(\frac{45 x^3}{7}\right) \left(\frac{3 x^3}{2}\right)=\frac{135 x^6}{14} \]

OpenStudy (anonymous):

wow hehe XD

OpenStudy (anonymous):

so who gets the medal reward?

OpenStudy (anonymous):

no one :P jk

OpenStudy (anonymous):

dj answered first